OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x+ 4.

OpenStudy (anonymous):

Hint(1)=To find the equation of a line Use the law y2-y1=m(x2-x1) . Hint(2)=when a Line is perpendicular to another line then Slope of the first line times the slope of the other equal -1 , M1*M2=-1 You have m2 which is -1/2 subs. ^ and get the another slope by solving the equation and then fill up the first rule ... ________________ Can you go through it now ?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Good ,If u need any help ,Inform me .

OpenStudy (anonymous):

I got it, thank you!

OpenStudy (anonymous):

Yw.